Use the following demand schedule on pens to answer the question.


Price Quantity

19 10 000

18 15 000

17 18 000


 Elas table pens What is the price elasticity of demand for pens in the price range R19 to R17? 



1
Expert's answer
2021-07-14T13:47:58-0400

"Pe=\frac{(\frac{Qd2-qd1}{q1})}{(\frac{p2-p1}{p1})}"

"=\frac{(\frac{18000-10000}{10000})}{(\frac{17-19}{19})}"


"=\frac{\frac{8000}{10000}}{\frac{2}{19}}"

"=\frac{0.8}{0.222222}\\=-7.6"